OpenStudy (anonymous):

The high school marching band has a total of 100 students. There are 3 times as many girls in the band as boys. How many girls are in the band?

OpenStudy (anonymous):

75

OpenStudy (anonymous):

Thanks @bluemustangjosh

OpenStudy (pfenn1):

y=no of girls in band x=no of boys in band y=3x x+y=100

OpenStudy (albert0898):

100 Students. Girls = Boys * 3 Boys + Girls = 4 100 / 4 = 25 25 * 3 = 75 75 Girls SHOW WORK :)

OpenStudy (anonymous):

Thankks

OpenStudy (albert0898):

Anytime!